On average, 3 customers make purchases every 2 hours at a certain hardware. Assume that the number of customers making purchases follows a Poisson distribution. If a customer has just made a purchase, calculate (to 4 decimal places) the probability that the time until the next customer arrives to make a purchase is less than 7 minutes:


1
Expert's answer
2021-11-15T02:39:05-0500

"\begin{aligned}

\lambda &=\frac{3}{2 \times 60} \\

&=3 / 120 \\

\lambda &=0.025

\end{aligned}"

By using Poisson distribution "P(X=x)=\frac{e^{-\lambda} \cdot \lambda^{x}}{x!}" ,

"\begin{aligned}

P(X<7) &=P(X=0)+P(X=1)+P(X=2)+P(X=3)+P(X=4)+P(X=5)+P(X=6) \\

&=0.95731+0.02438+0.00330+0.0000+00000+0.000+0.0000 \\

P(X<7) &=1.0000

\end{aligned}"